# Ctrl-GenAug: Controllable Generative Augmentation for Medical Sequence Classification

### v1

@@ III-C Multimodal Conditions Guidance
We consider four multimodal conditions to ensure comprehensive and accurate control over the sequence synthesis procedure (see Figs. 2-3). Specifically, these conditions are divided into semantic and sequential ones for visual appearance control and serial guidance, respectively. We highlight that our generator supports composable synthesis by allowing users to flexibly choose any single condition or composition during inference. This flexibility makes our generator particularly user-friendly, as it enables high-quality synthesis even when certain conditions are missing. Details are described below.
III-C1 Semantic Conditions for Visual Appearance Control
@@
- Descriptive Text: It provides an intuitive indication of the coarse-grained semantic concepts of sequences. In our implementation, we pre-align the visual and textual features by finetuning the widely used CLIP model [59] using our medical datasets to mitigate the large gap between them. Then, semantic information of texts can feasibly query relevant medical visual patterns, thus easing subsequent training. Similar to [8], the text guides the generator via cross-attention layers.
- Class Label: Generating medical data with the expected disease class is crucial for enhancing downstream classification, particularly improving the diagnostic accuracy of underrepresented high-risk sets. Thus, we propose encouraging the model to focus on the grading signal. Specifically, rather than merging the class label with descriptive text, we treat it as a separate tag to directly provide disease-specific guidance, as merging with text may dilute the class signal. We further propose a dedicated insertion method for this purpose. First, similar to the timestep embedding process [8] in diffusion models, we adopt a class label encoder composed of a label discretization layer and an embedding layer to obtain its embeddings. The addition between the timestep and class label embeddings is then input to the spatial convolution layer in each UNet block for visual appearance control.
- Image Prior: Merely using the above conditions faces challenges of insufficient semantic control and an inevitable domain gap between real and synthetic sequences, constraining the capability of generative augmentation. Hence, we introduce the first frame/slice of real-domain sequences as image priors to provide rich semantic guidance and yield real-domain style sequences. As image prior offers global semantics like texts, we align image prior features $f_{i}$ extracted from image prior encoder (see Fig. 3) with text embeddings $f_{t}$ generated by our medical data-specific CLIP text encoder for joint guidance. Specifically, motivated by [60], we replace the text-guided cross-attention layers in original UNet blocks with decoupled counterparts that handle texts and image priors in parallel and then merge the results by addition. The decoupled attention can be formulated as:
@@
III-C2 Sequential Condition for Serial Guidance
Apart from perceiving visual concepts, modeling sequential knowledge is also important in sequence synthesis, with sequential cues playing a vital role in this process. Most existing methods [35, 36] adopted dense optical flow [34] to promote sequential dynamics modeling. However, optical flow extraction requires high computational demands [33]. Thus, we instead introduce computation-efficient motion field as a sequential condition in sequence LDM. It explicitly showcases the pixel-wise motions between adjacent frames/slices (see Figs. 2-3). In our approach, we first extract the motion fields of real sequences using a Python package [61]. Then, a motion encoder (see Fig. 3) receives the motion fields and produces motion features. Last, the features are concatenated with latent representations $\mathit{z}$ from VAE along the channel dimension for serial guidance.
@@ III-C2 Sequential Condition for Serial Guidance
III-D Sequential Augmentation Module
@@
III-D1 Key-frame/slice Attention
@@
III-D2 Motion Field Attention
To further boost the sequential coherence of the generated sequences, we propose to reuse the motion field condition (refer to Sec. III-C2) by introducing a motion field attention (MFA) mechanism after KA. As shown in Fig. 4(c), MFA requires the patches to communicate with those in the same motion field-based pathway including itself, which eliminates flickers of the generated sequences to make the contents visually smooth. Inspired by [62], MFA is implemented in two steps: a) motion field-based patch pathway sampling and b) attention calculation. In step 1, we sample the patch pathways based on the m-scaled downsampled motion fields. Take a video sequence as an example, for a patch $p_{l}$ on the l-th frame of a f-frame video, the $path$ can be derived from the motion field. Since the sampling procedure inevitably generates multiple pathways for the same patch, we randomly sample a pathway to ensure its uniqueness to which each patch belongs. In this setting, let $H,W$ the height and width of the input video frame, then the size of the pathway set after sampling equals $\frac{H\times W}{m^{2}}\times f$. In step 2, we calculate $Attention(\mathbf{Q},\mathbf{K},\mathbf{V})$ with:
@@ III-D2 Motion Field Attention
III-E Noisy Synthetic Data Filter
@@
We use the proposed sequence generator to constitute the synthetic samples set. Concretely, assuming there are $n$ training clips in the target dataset, with a bank of conditions derived from each clip, we synthesize a group of clips guided by each conditions bank. Eventually, we obtain $n$ groups of clips to form our synthetic sample set, with a total of $N$ clips. Although visually realistic and smooth (good cases, Fig. 5), sequence synthesis may still suffer from class semantics misalignment, cross-frame/slice inconsistency or over-consistency (i.e., almost static clip), and inter-clip similarity. For instance, in Fig. 5(e), the synthesized carotid clip category is wrong, which should be moderate rather than mild. In Fig. 5(f), the generated clip includes abrupt changes in anatomical structures. Hence, blindly using all synthetic clips for classifier learning will significantly cause a performance drop due to noisy samples. Our work proposes a noisy synthetic data filter to adaptively remove harmful generated clips at class semantics and sequential levels, as illustrated in Algorithm 1. This can also link the synthetic data to downstream tasks, thus potentially achieving a higher performance upper bound.
III-E1 Class Semantics Misalignment Filtering
@@
This scheme is designed to filter the generated clips whose visual contents mismatch the corresponding class labels. Motivated by anomaly detection [63], we consider that such noisy synthetic samples will produce large losses if fed into a powerful classifier trained on real ones. Hence, a loss threshold can be set to identify and filter the anomalistic clips. Specifically, we first adopt a well-trained classifier in the real data domain to calculate case-wise cross-entropy loss values $\left\{l_{x}\right\}_{x=1}^{M}$ based on corresponding category $c$ for all synthetic clips in each group. Then, we set the average loss as the noise identification threshold of $c$. Last, we conduct class semantics misalignment filtering at the group level. If the loss of a synthetic clip with $c$ exceeds the threshold $L_{c}$, it will be viewed as a noisy sample and excluded from the synthetic pool.
III-E2 Sequential Filtering
@@
This strategy aims to filter noisy samples at a) inner-sequence and b) inter-sequence levels. For a), we screen out synthetic clips with gentle dynamic, avoiding those that are either inconsistent or over-consistent affecting downstream learning. Specifically, we retain a synthetic clip whose cross-frame/slice consistency falls in a pre-computed range. It is determined by K-means clustering [64] based on the cross-frame/slice consistency of all synthetic clips. To assess the video sequence consistency, [33] tends to use CLIP [59] image embeddings to compute the average cosine similarity across consecutive frames. However, this method leads to inadequate evaluation due to the usage of limited informative CLIP embeddings. To resolve the problem, we propose a metric termed VAE-Seq that utilizes latent embeddings from pretrained VAE, instead of deriving embeddings from CLIP space, to assess cross-frame/slice consistency. Therefore, with the higher dimension of latents (e.g., 4$\times$32$\times$32 for input size 256$\times$256 compared to 768 in CLIP embeddings), VAE-Seq reflects an accurate evaluation using more fine-grained and informative features. For b), we seek to diversify the samples in the synthetic set to prevent overfitting and avoid wasting computational resources in downstream training. To this end, we perform the inter-sequence filtering based on inter-clip similarity in each group, as shown in Algorithm 1. We quantify the inter-clip similarity by calculating the cosine distance of VAE latent embeddings between frame/slice pairs in two clips and averaging the obtained distance values.
III-F Evaluation
@@
As a common practice, existing sequence synthesis works evaluate the synthetic quality by mostly using Fréchet Inception Distance (FID) [65] and Fréchet Video Distance (FVD) [66]. However, researchers have verified that these metrics do not consistently correlate with performance metrics on downstream tasks [46, 7]. To exhaustively evaluate the values of synthetic samples to downstream tasks, eleven popular classifiers are deployed for comparison with respect to diagnostic accuracy and area under the receiver operating characteristic curve (AUROC). Besides, we adopt the proposed VAE-Seq to assess the cross-frame/slice consistency (refer to Sec. III-E2). Moreover, we consider it crucial to evaluate the smoothness of generated clips. Motivated by [67], we quantify it using the proposed metric (termed Dynamic Smoothness), which is the mean absolute error between the reconstructed frames/slices from the interpolation model [68] and the original real ones.
@@
IV Experiments
In this section, we first comprehensively validated the effectiveness and generality of our method on three medical datasets, using eleven popular networks trained on three paradigms. We then tested the performance of our method in underrepresented high-risk sets and out-domain conditions. Last, we analyzed the contributions of our proposed multimodal conditions guidance, sequential augmentation module, and noisy synthetic data filter respectively, as well as the impacts of traditional and generative augmentations.
TABLE I: Experimental datasets and settings.
Carotid TUSC [69] ACDC [70]
@@ IV Experiments
Classifier-free guidance [71] factor(7.5)   Sampling steps(200)
@@
IV-A Datasets and Settings
Carotid dataset was collected by three medical centers with approval from local institutional review boards, including a) The Third Affiliated Hospital of Sun Yat-sen University, b) Shenzhen Longgang District People’s Hospital, and c) The Eighth Affiliated Hospital of Sun Yat-sen University. It consists of 273 US video sequences from 231 patients with carotid plaque-induced stenosis. The severity of carotid stenosis is graded into three stages, i.e., mild, moderate, and severe. We uniformly sampled several non-overlapping 8-frame clips from each sequence. The final dataset involved an in-domain (ID) subset (407 clips from 193 patients collected by Center (a)) and an out-domain (OD) subset (79 clips from 38 patients acquired from Center (b)-(c)). The ID/OD subset was randomly split into 309/39 and 98/40 clips for training and testing at the patient level. Each clip has up to two plaques and was labeled with stenosis grading and descriptive text by experienced sonographers using Pair annotation software package [72]. The text annotations indicated the plaque characteristics of echogenicity and location (see Fig. 2). We cropped the region of interest based on the pretrained carotid vessel detector [73] for easing the model learning.
TUSC [69] contains 192 thyroid US video sequences from 167 patients collected by the Stanford University Medical Center, with each including one nodule. Each sequence provided a TI-RADS level and nodule descriptors [74] (e.g., shape), the latter of which was used to form descriptive text. The TI-RADS level is ranging from 1 to 5. Considering the imbalanced distribution of TI-RADS levels, as suggested by the experienced sonographers, we rebuilt the dataset by removing the TI-RADS1 video sequence and combining the levels 2-3 sequences into a unified category named TI-RADS2-3. We uniformly sampled the sequences, obtaining 633 8-frame clips, with 448 for training and 185 for testing. Each sampled clip inherited the same TI-RADS level and text as the original sequence. The bounding box of the nodule region for each frame was extracted by radiologists, similar to [75].
ACDC [70] includes 4D cardiac cine-MRI from 150 patients acquired at the University Hospital of Dijon. The patients were evenly divided into five classes: normal patients (NOR), myocardial infarction (MINF), dilated cardiomyopathy (DCM), hypertrophic cardiomyopathy (HCM), and abnormal right ventricle (ARV). We first extracted 3D volume sequences at end-systole and diastole phases from each case and assigned corresponding classes. Then, the 8-slice clips were uniformly sampled from each sequence along the slice axis. Finally, 205 and 109 clips were constructed for training and testing. Text prompts were set to empty strings during generator training.
Table I presents a detailed description of experimental datasets and settings. Using all available conditions, we synthesized 300 clips per category for the ACDC dataset and 500 clips per class for others. All frames/slices were resized to 256$\times$256. By default, we set m=8 for the MFA mechanism. For the sequence generator, the pretraining LDM was built upon Stable Diffusion [8] and initialized with the public weights [^2] https://huggingface.co/CompVis/stable-diffusion-v1-4. During sequence LDM finetuning, we updated the parameters of SAM, SA layers, and motion encoder. We further refined the text- and image prior-sequence alignment by keeping the query projection matrices in cross-attention layers trainable. During generator training, we employed the multimodal conditions joint training strategy [76]. This allows users to flexibly select any single condition or combination during sampling, without finetuning the model for each specific combination. Motivated by [46], we adopted three paradigms to explore the impacts of the synthetic samples for downstream classifiers. Baseline: merely trained with real clips. Real-finetune: initially pretrained with synthetic clips and then finetuned with real ones. Joint-train: real clips were first over-sampled to match the number of synthetic clips, the classifier was then trained on the real-synthetic mixed set. All classification experiments were conducted on MMAction2 [77] except for FTC [78] and CSG-3DCT [1].
@@ IV-A Datasets and Settings
Method Backbone Training Paradigm Carotid TUSC [69] ACDC [70]
@@
IV-B Generative Augmentation-based Sequence Classification

### v2

@@ 3.3 Multimodal Conditions Guidance
We consider four multimodal conditions to ensure comprehensive and accurate control over the sequence synthesis procedure (see Figs. 3-4). Specifically, these conditions are divided into semantic and sequential ones for visual appearance control and serial guidance, respectively. We highlight that our generator supports composable synthesis by allowing users to flexibly choose any single condition or composition during inference. This flexibility makes our generator particularly user-friendly, as it enables high-quality synthesis even when certain conditions are missing. Details are described below.
3.3.1 Semantic Conditions for Visual Appearance Control
@@
- Descriptive Text: It provides an intuitive indication of the coarse-grained semantic concepts of sequences. In our implementation, we pre-align the visual and textual features by finetuning the widely used CLIP model (Radford et al., 2021) using our medical datasets to mitigate the large gap between them. Then, semantic information of texts can feasibly query relevant medical visual patterns, thus easing subsequent training. Similar to (Rombach et al., 2022), the text guides the generator via cross-attention layers.
- Class Label: Generating medical data with the expected disease class is crucial for enhancing downstream classification, particularly improving the diagnostic accuracy of underrepresented high-risk sets. Thus, we propose encouraging the model to focus on the grading signal. Specifically, rather than merging the class label with descriptive text, we treat it as a separate tag to directly provide disease-specific guidance, as merging with text may dilute the class signal. We further propose a dedicated insertion method for this purpose. First, similar to the timestep embedding process (Rombach et al., 2022) in diffusion models, we adopt a class label encoder composed of a label discretization layer and an embedding layer to obtain its embeddings. The addition between the timestep and class label embeddings is then input to the spatial convolution layer in each UNet block for visual appearance control.
- Image Prior: Merely using the above conditions faces challenges of insufficient semantic control and an inevitable domain gap between real and synthetic sequences, constraining the capability of generative augmentation. Hence, we introduce the first frame/slice of real-domain sequences as image priors to provide rich semantic guidance and yield real-domain style sequences. As image prior offers global semantics like texts, we align image prior features $f_{i}$ extracted from image prior encoder (see Fig. 4) with text embeddings $f_{t}$ generated by our medical data-specific CLIP text encoder for joint guidance. Specifically, motivated by (Ye et al., 2023), we replace the text-guided cross-attention layers in original UNet blocks with decoupled counterparts that handle texts and image priors in parallel and then merge the results by addition. The decoupled attention can be formulated as:
@@
3.3.2 Sequential Condition for Serial Guidance
Apart from perceiving visual concepts, modeling sequential knowledge is also important in sequence synthesis, with sequential cues playing a vital role. Most existing methods (Liang et al., 2023; Ni et al., 2023) adopted dense optical flow (Zach et al., 2007) to promote sequential dynamics modeling. However, optical flow extraction requires high computational demands (Wang et al., 2024b). Thus, we instead introduce computation-efficient motion field as a sequential condition in sequence LDM. It explicitly showcases the pixel-wise motions between adjacent frames/slices (Figs. 3-4). In our approach, we first extract the motion fields of real sequences using a Python package (Bommes et al., 2020). Then, a motion encoder (Fig. 4) receives the motion fields and produces motion features. Last, the features are concatenated with latent representations $\mathit{z}$ from VAE along the channel dimension for serial guidance.
@@ 3.3.2 Sequential Condition for Serial Guidance
3.4 Sequential Augmentation Module
@@
3.4.1 Key-frame/slice Attention
@@
3.4.2 Motion Field Attention
To further boost the sequential coherence of the generated sequences, we propose to reuse the motion field condition (refer to Sec. 3.3.2) by introducing a motion field attention (MFA) mechanism after KA. As shown in Fig. 5(c), MFA requires the patches to communicate with those in the same motion field-based pathway including itself, which eliminates flickers of the generated sequences to make the contents visually smooth. Inspired by (Cong et al., 2024), MFA is implemented in two steps: a) motion field-based patch pathway sampling and b) attention calculation. In step 1, we sample the patch pathways based on the m-scaled downsampled motion fields. Take a video sequence as an example, for a patch $p_{l}$ on the l-th frame of a f-frame video, the $path$ can be derived from the motion field. Since the sampling procedure inevitably generates multiple pathways for the same patch, we randomly sample a pathway to ensure its uniqueness to which each patch belongs. In this setting, let $H,W$ the height and width of the input video frame, then the size of the pathway set after sampling equals $\frac{H\times W}{m^{2}}\times f$. In step 2, we calculate $Attention(\mathbf{Q},\mathbf{K},\mathbf{V})$ with:
@@ 3.4.2 Motion Field Attention
3.5 Noisy Synthetic Data Filter
We use the proposed sequence generator to constitute the synthetic samples set. Concretely, assuming there are $n$ training clips in the target dataset, with a bank of conditions derived from each clip, we synthesize a group of clips guided by each conditions bank. Eventually, we obtain $n$ groups of clips to form our synthetic sample set, with a total of $N$ clips.
@@ 3.5 Noisy Synthetic Data Filter
Although visually realistic and smooth (good cases, Fig. 6), sequence synthesis may still suffer from class semantics misalignment, cross-frame/slice inconsistency or over-consistency (i.e., almost static clip), and inter-clip similarity. For instance, in Fig. 6(e), the synthesized carotid clip category is wrong, which should be moderate rather than mild. In Fig. 6(f), the generated clip includes abrupt changes in anatomical structures. Hence, blindly using all synthetic clips for classifier learning will significantly cause a performance drop due to noisy samples. Our work proposes a noisy synthetic data filter to adaptively remove harmful generated clips at class semantics and sequential levels, as illustrated in Algorithm 1. This can also link the synthetic data to downstream tasks, thus potentially achieving a higher performance upper bound.
3.5.1 Class Semantics Misalignment Filtering
@@
This scheme is designed to filter the generated clips whose visual contents mismatch the corresponding class labels. Motivated by anomaly detection (Kwon et al., 2020), we consider that such noisy synthetic samples will produce large losses if fed into a powerful classifier trained on real ones. Hence, a loss threshold can be set to identify and filter the anomalistic clips. Specifically, we first adopt a well-trained classifier in the real data domain to calculate case-wise cross-entropy loss values $\left\{l_{x}\right\}_{x=1}^{M}$ based on corresponding category $c$ for all synthetic clips in each group. Then, we set the average loss as the noise identification threshold of $c$. Last, we conduct class semantics misalignment filtering at the group level. If the loss of a synthetic clip with $c$ exceeds the threshold $L_{c}$, it will be viewed as a noisy sample and excluded from the synthetic pool.
3.5.2 Sequential Filtering
@@
This strategy aims to filter noisy samples at a) inner-sequence and b) inter-sequence levels. For a), we screen out synthetic clips with gentle dynamic, avoiding those that are either inconsistent or over-consistent affecting downstream learning. Specifically, we retain a synthetic clip whose cross-frame/slice consistency falls in a pre-computed range. It is determined by K-means clustering (Macqueen, 1967) based on the cross-frame/slice consistency of all synthetic clips. To assess the video sequence consistency, (Wang et al., 2024b) tends to use CLIP (Radford et al., 2021) image embeddings to compute the average cosine similarity across consecutive frames. However, this method leads to inadequate evaluation due to the usage of limited informative CLIP embeddings. To resolve the problem, we propose a metric termed VAE-Seq that utilizes latent embeddings from pretrained VAE, instead of deriving embeddings from CLIP space, to assess cross-frame/slice consistency. Therefore, with the higher dimension of latents (e.g., 4$\times$32$\times$32 for input size 256$\times$256 compared to 768 in CLIP embeddings), VAE-Seq reflects an accurate evaluation using more fine-grained and informative features. For b), we seek to diversify the samples in the synthetic set to prevent overfitting and avoid wasting computational resources in downstream training. To this end, we perform the inter-sequence filtering based on inter-clip similarity in each group, as shown in Algorithm 1. We quantify the inter-clip similarity by calculating the cosine distance of VAE latent embeddings between frame/slice pairs in two clips and averaging the obtained distance values.
3.6 Evaluation
@@
Existing sequence synthesis works evaluate the synthetic quality by mostly using Fréchet Inception Distance (FID) (Heusel et al., 2017) and Fréchet Video Distance (FVD) (Unterthiner et al., 2018). However, researchers have verified that these metrics do not consistently correlate with performance metrics on downstream tasks (Yang et al., 2024; Luo et al., 2024). To exhaustively evaluate the values of synthetic cases to downstream tasks, 11 popular classifiers are deployed for comparison with respect to diagnostic accuracy and area under the receiver operating characteristic curve (AUROC). Besides, we adopt the proposed VAE-Seq to assess the cross-frame/slice consistency (more details refer to Sec. 3.5.2). Moreover, motivated by (Huang et al., 2024c), we quantify the smoothness of generated clips using the metric termed Dynamic Smoothness, which is the mean absolute error between the reconstructed frames/slices from the interpolation model (Li et al., 2023) and the original real ones.
@@
4 Experiments
We first comprehensively validated the effectiveness and generality of our Ctrl-GenAug on 3 medical datasets, using 11 popular networks trained on 3 paradigms. We then tested the performance of Ctrl-GenAug in underrepresented high-risk sets and out-domain conditions. Last, we analyzed the contributions of our multimodal conditions guidance, sequential augmentation module, and noisy synthetic data filter, respectively; as well as the impacts of traditional and generative augmentations.
4.1 Datasets and Settings
Carotid was collected by three medical centers with approval from local institutional review boards, including a) The Third Affiliated Hospital of Sun Yat-sen University, b) Shenzhen Longgang District People’s Hospital, and c) The Eighth Affiliated Hospital of Sun Yat-sen University. It consists of 273 US video sequences from 231 patients with carotid plaque-induced stenosis. The severity of carotid stenosis is graded into three stages, i.e., mild, moderate, and severe. We uniformly sampled several non-overlapping 8-frame clips from each sequence. The final dataset involved an in-domain (ID) subset (407 clips from 193 patients collected by Center (a)) and an out-domain (OD) subset (79 clips from 38 patients acquired from Center (b)-(c)). The ID/OD subset was randomly split into 309/39 and 98/40 clips for training and testing at the patient level. Each clip has up to two plaques and was labeled with stenosis grading and descriptive text by experienced sonographers using Pair annotation software package (Liang et al., 2022). The text annotations indicated the plaque characteristics of echogenicity and location (see Fig. 3). We cropped the region of interest based on the pretrained carotid vessel detector (Liu et al., 2023) for easing the model learning.
TUSC (Stanford-AIMI, 2021) contains 192 thyroid US video sequences from 167 patients collected by the Stanford University Medical Center, with each including one nodule. Each sequence provided a TI-RADS level and nodule descriptors (Tessler et al., 2017) (e.g., shape), the latter of which was used to form descriptive text. The TI-RADS level is ranging from 1 to 5. Considering the imbalanced distribution of TI-RADS levels, as suggested by the experienced sonographers, we rebuilt the dataset by removing the TI-RADS1 video sequence and combining the levels 2-3 sequences into a unified category named TI-RADS2-3. We uniformly sampled the sequences, obtaining 633 8-frame clips, with 448 for training and 185 for testing. Each sampled clip inherited the same TI-RADS level and text as the original sequence. The bounding box of the nodule region for each frame was extracted by radiologists, similar to (Yamashita et al., 2022).
ACDC (Bernard et al., 2018) includes 4D cardiac cine-MRI from 150 patients acquired at the University Hospital of Dijon. They were evenly divided into five classes: normal patients (NOR), myocardial infarction (MINF), dilated cardiomyopathy (DCM), hypertrophic cardiomyopathy (HCM), and abnormal right ventricle (ARV). We first extracted volume sequences at end-systole and diastole phases from each case and assigned corresponding classes. Then, 8-slice clips were uniformly sampled from each sequence along the slice axis. Finally, 205 and 109 clips were obtained for training and testing. Text prompts were set to empty strings during generator training.
Table 1 presents a detailed description of experimental datasets and settings. Using all available conditions, we synthesized 300 clips per category for the ACDC dataset and 500 clips per class for others. All frames/slices were resized to 256$\times$256. By default, we set m=8 for the MFA mechanism. For the sequence generator, the pretraining LDM was built upon Stable Diffusion (Rombach et al., 2022) and initialized with the public weights [^4] https://huggingface.co/CompVis/stable-diffusion-v1-4. During sequence LDM finetuning, we updated the parameters of SAM, SA layers, and motion encoder. We further refined the text- and image prior-sequence alignment by keeping the query projection matrices in cross-attention layers trainable. During generator training, we employed the multimodal conditions joint training strategy (Huang et al., 2023a). This allows users to flexibly select any single condition or combination during sampling, without finetuning the model for each specific combination. Motivated by (Yang et al., 2024), we adopted three paradigms to explore the impacts of the synthetic samples for downstream classifiers. Baseline: merely trained with real clips. Real-finetune: initially pretrained with synthetic clips and then finetuned with real ones. Joint-train: real clips were first over-sampled to match the number of synthetic clips, the classifier was then trained on the real-synthetic mixed set. All classification experiments were conducted on MMAction2 (Contributors, 2020) except for FTC (Ahmadi et al., 2023) and CSG-3DCT (Zhou et al., 2023).
Table 1: Experimental datasets and settings. CF, Classifier-free.
Carotid TUSC ACDC
@@ 4.1 Datasets and Settings
CF guidance (Ho and Salimans, 2022) factor(7.5)   Sampling steps(200)
@@
Method Backbone Training Paradigm Carotid TUSC (Stanford-AIMI, 2021) ACDC (Bernard et al., 2018)
@@
4.2 Generative Augmentation-based Sequence Classification